a 1000.kg car moving at north at 100. km/hr brakes to a stop in 50. m within a uniform acceleration. what are magnitude and direction of the net force on the car ?
1
Expert's answer
2019-10-31T11:14:27-0400
"F=ma"

"a=\frac{v^2}{2s}"

"F=m\frac{v^2}{2s}=1000\frac{(\frac{100}{3.6})^2}{2(50)}=7716\ N"

Direction: at south.